The probabilities that players A and B of a team are selected for the captaincy for a tournament are 0.6 and 0.4, respectively. If A is selected the captain, the probability that the team wins the tournament is 0.8 and if B is selected the captain, the probability that the team wins the tournament is 0.7. Then the probability, that the team wins the tournament, is :
- 0.74
- 0.76
- 0.72
- 0.78
Answer
(2) 0.76
Solution
Let E1: A is the captain, E2: B is the captain, W: the team wins.
P(E1)=0.6, P(E2)=0.4, P(W E1)=0.8, P(W E2)=0.7
E1, E2 are mutually exclusive and exhaustive, by the theorem of total probability:
P(W)=P(E1)P(W E1)+P(E2)P(W E2)
P(W)=0.6 0.8+0.4 0.7=0.48+0.28=0.76
Answer: (2)